OpenStudy (anonymous):

Find the area of a triangle whose base is 2√3 in. and height is 3√3 in. 9 in.² 18 in.²

OpenStudy (studygurl14):

\(\Large\frac{2\sqrt{3}(3\sqrt{3})}{2}\tt =~the~answer\)

OpenStudy (anonymous):

9?

OpenStudy (studygurl14):

\(\Large\tt\color{blue}{CORRECT!}\)

OpenStudy (anonymous):

thnx:)

OpenStudy (studygurl14):

anytime \(\Large\ddot\smile\)